a retailer incurs a fixed cost of N$330 when purchasing sugar for his stock. he pays N$15.00per packet which he resell at N$18.00per packet.how many packets should he purchase and sell in order to break even?
Answer by mananth(16946)

cost = sale ( break even)
let him by x packets
5x+330 = 18x
18x-15x=330
3x=330
/3
x=110 packets